1. Prepare stock solution of 10mM Fe3+ at a low pH. ( Mix 1.62g of FeCl3 and 2ml HCL stock solution in a volumetric flask and fill up to 1L with distilled water. )

2. Dilute Fe3+ stock solution 10times, add 50ml of them into an Erlenmeyer flask and after inserting it into the heating mantle, wait for the constant temperature condition to complete.

3. How many grams of H2O2 should be added into 50ml of 10 times diluted Fe3+ stock solution so that the 34.5wt% H2O2 stock solution turns into a H2O2 solution with 100mM concentration?
Answer #1

1, the first solution is 10mM as 1,62 gm is made upto 1000ml. ( as mol wt of ferric chloride is 162 gm)

then dilute it 10times means 1mM solution .  

then take 50ml solution of the above solution means 50 ml X 1mM Ferric chloride solution.

the millimoles of ferrice solution is 50mM.

the reaction of peroxide H2O2 and ferric chloride equation is as given below.

peroxide is dissociated to give water and oxygen and ferric ion becomes ferrous.

since Fe 3+ - to   Fe 2+  + e

and H2O2 gives rise to H2O and O2 but in electronic equation

H2O2 reacts to form H2O and O + 2e

2 moilecule ferricis reacted with one molecule H2O2

Hydrogen Peroxide is H2O2 with a molar mass of 34.015 g/mole this is done by (H =1 and O =16 )

A 34.5% solution of H2O2 is 34.5% H2O2 and 65.5% H2O by volume.

So a liter of this solution contains 0.345 liter of H2O2 and 0.65 liter

of H2O.

The density of H2O2 is 1.44 g/cm^3 = 1.44 g / 0.001 liter.

0.345 liter of H2O2 has a mass of (0.345 liter)(1.44 g / 0.001 liter) =

496.8  g, which is 496,8 g / 34.015 g/mole = 14.611  moles of H2O2.

This means 34.5 % H2O2 solution = 14.611 M H2O2. or 14611 mM H2O2.

Now to get 100mM solution after neutralizing 50mM of ferric chloride solution that means

total 100mM H2O2 and 50/2 = 25 mM total 125mM SOLUTION WAS AT THE BEGINNING.

Now one litre 34.5% H2O2 means 14611 mM solution

or 14611 mM peroxide solution means 1000 ml or 1440gm solution

so 125 mM ............................................ (125X1000 ) /14611 ml or 8.555  ml or

1,44 X 8.555 gm = 12.319 gm.

So answer is 12.319 gm of peroxide should be added.